You are tasked with estimating the cost of capital for a firm. The risk-free rate is 4.7%, the expected rate of return on the market is 10.7%. Now, another similar company (similar unlevered cost of capital) has a debt-to-equity ratio of 1 to 2. It has a debt beta near zero and an equity market-beta of 1.5. Your own firm has more debt, for a debt-to-equity ratio of 1 to 1, with a debt beta of 0.3. What is a good estimate for your firm's cost of capital (WACC)?
Answer #1

Given,

Risk free rate Rf = 4.7%

expected rate of return on the market Rm = 10.7%

For other firm, it is given that

D/E ratio = 1:2

levered beta =  equity market-beta = 1.5

So unlevered beta = Levered beta (equity beta)​/(1 + (1-tax rate)*(D/E))

Assuming tax rate = 0

So, unlevered beta = 1.5/(1 + 1/2) = 1

For our firm,

D/E ratio = 1:1

unlevered beta = 1 (from the other firm)

So, levered beta = Unlevered beta*(1 + (1-tax rate)*(D/E)) = 1*(1 + 1) = 2

So, cost of equity Ke of the firm using CAPM = Rf + equity beta*(Rm - Rf) = 4.7 + 2*(10.7-4.7) = 16.7%

cost of debt Kd using same = Rf + debt beta*(Rm-Rf) = 4.7 + 0.3*(10.7-4.7) = 6.5%

So, WACC = Wd*Kd + We*Ke = 0.5 * 6.5 + 0.5 * 16.7 = 11.6%